OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (anonymous):

So far this is what I've got... <2,-4> x <3,-8> = 6+32= 38 2^2 = 4 -4^2= -16 4+-16= -12 sqrt 12 3^2 = 9 -8^2 = -64 9+(-64)= -55 sqrt55 And then I did cos-1 (38/sqrt12 sqrt55) and then that gives me error so I don't know what I'm doing wrong...

OpenStudy (anonymous):



OpenStudy (anonymous):

But I got error so where did I mess up?